Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: In A configuration (08 Jun). Holiday next Monday (so this meeting will likely start late). New versions of PCMC; CBE last week; peculiarities with BlB behavior; CBE issues with MPI (missing scans not completely gone - new guidelines to avoid). Call for Proposals goes out on Thursday. Discussions today on scan alignment, data rates, and an autoflagging demo.
  • Key dates:
    • 30 Jun - Call for Proposals
    • 08 Jun - 12 Sep: A configuration (move from 31 May - 08 Jun).
    • 01 Aug - Proposal Deadline (C, CnB, B)
  • Staff Travel:
    • LS: 9 May - 20 Jun
    • GvM: 25 May - 24 Jun
    • CC: 19 Jun - 26 Jun
    • MC: 18 Jun - 26 Jun
    • CC: 28 Jun - 30 Jun (CARMA TAC)
    • RP: 10 Jul - 23 Jul (UK, SA)
    • RP: 09 Aug - 21 Aug (URSI)
    • FO: 12 Aug - 12 Sep
    • RP: 10 Sep - 16 Sep (Brazil)
  • Visitors:
    • Robert Harris (AC982): 24 Apr - 16 Jul (High frequency w/ Claire)
    • Michiel Brentjens: 11 Jun - 11 Nov (Holography; characterizing polarized beams)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Antennas: EA21 back to array 11 Jul. ACU meeting Wednesday at 1530.
    • Receivers: EA12 X band; EA21 spate of receivers by the end of the month (28 Jun)
    • Calibration (Perley): Gain compression/expansion still being looked at.
  • WIDAR (Rupen):
    • Notes (Rupen/Sowinski): New PCMC last week; currently have no spares in either BlB/StB; working to rectify. Issues over the weekend with BlBs (both classic effing problem and the new partial shutdown which ultimately leads to a loss of communication.
  • Computing (Robnett):

Software status and updates (Butler)

  • Links:
  • SSS: Work on new OSRO/RSRO capabilities. Correlator: CBE work to hunt down issues with missing scans from the weekend. Work to push on data rates.
  • Data Rate Discussion: Currently promising 25 MB/s; we're capable of approximately 15 MB/s. Work to improve warnings and notification about the limitations.
  • CASA Update (new functionality available/issues/enhancements):Notes: new Autoflagging capabilities.
  • AIPS Update (new functionality available/issues/enhancements): Notes: Eric out today; more next week.
  • SDM Update: Notes: Ephemeris table status.

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler)

Configuration move week (work around these activities:
  • Mon
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Tue
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Thu:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Fri (not-yet-a-holiday):
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Sat-Sun
    • Science operations + planned testing

Issues Affecting Science Observing:

  • Correlator: [Rupen/Carlson w/ Momjian/Sault] V stokes issues.Further data taking; new theories; Bob leaves on Thursday; discussion on Tuesday.
  • Correlator: [Pokorny/Rupen/McMullin] time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags).
    • Report (Rupen):
  • Correlator: [Clark/Carlson] Delay rocking across bands. Ongoing discussion of implementation; pending Bryan's return
  • Array: [Gustaaf] Amplitude wobbles seen on subset of antennas; phase jumps at scan boundaries also seen on subset of antennas (Ka, Q bands only thus far).
  • Array: [All] Zoo of other effects (some data are very good, some still seeing antenna-based IF/Polarization issue).

Testing This week/soon

  • CM/CBE: Data rate testing and push; work on new missing scans.
  • V stokes testing (Momjian/Greisen/Rupen): Further testing (using well understood (VLA) observations).
  • Astrometry (Bruthaler): UVW precision tests.

  • Correlator: Rupen: S-band shadow testing
  • Spectral line: Momjian: HI High resolution bandpass testing
  • Array: *; Individual issue sleuthing (focus on phase jumps).

Commissioning Groups

  • Big deadline is end of this month for incorporation into links for Call for Proposals due 30 June.

    • 4-band (Owen):
      • Notes:
    • Adv: Low band (Owen):
      • Notes:
    • Adv: Astrometry (Mioduszewski):
      • Notes: Begin testing of UVW precision issues (Brunthaler).
  • Calibrator Models
    • Notes: A configuration data being collected.
  • HPC (Bhatnagar/Owen/Golap):
    • Notes:
  • Imaging (Ott/Golap):
    • Notes: SDM content of Mosaic modes discussion.
  • Algorithm (Bhatnagar, Owen):

RSRO/ECSO Reports

RSRO A configuration status: https://safe.nrao.edu/wiki/bin/view/EVLA/2011BRSROStatus

Individual reports (updates)

Misc:

  • RSRO Calendar: https://safe.nrao.edu/wiki/bin/view/EVLA/RSROCalendar
  • RSRO DnC, C config status: https://safe.nrao.edu/wiki/bin/view/EVLA/2010CRSROStatus
  • Test Project Codes == *TRSR*
  • Guidelines for RSRO observing: https://safe.nrao.edu/wiki/bin/view/EVLA/RSROObservingPreparationGuidelines
  • Notes:
    • Current Residents
      • D.J. Saikia (10C-119/AI145)::RSRO-75; 300G(2); x7106-POC: Momjian (Low Frequency)
      • J. Irwin (10C-119AI145)::RSRO-75; 300G(1); x7333 -- POC: Myers (Low Frequency)
      • Peter Hofner (10B-124/AH1011)::RSRO-26; 200G(1); x7496 -- POC: Ott (CASA testing)
      • Sergio Dzib (10B-131); 200G;x7496 -- POC: Mioduszewski (Astrometry)
      • Jacqueline van Gorkom (10C-211/AV324); 200D/x7098 -- POC: Momjian (Low Frequency); Not present
      • Laura Chomiuk (*); 200B/x7497 -- POC: Goss (Bandpass stability)
      • Andreas Brunthaler::200E/x7246 -- POC: Mioduszewski (Astrometry/VLBI)
      • Jane Greaves::(AC982)::RSRO-4; 300G;x7106 -- POC: Chandler (High Frequency Documentation for Workshop)

Demo Science Status

AOB

Misc Actions

  • Old/Low priority
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)

-- JosephMcMullin - 2011-06-27
Topic revision: r2 - 2011-06-27, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback